Block Treatment 1 2 3 4 Treatment Mean Tr1 2 1 2 3 2 Tr2 4...

Block
Treatment 1 2 3 4 Treatment Mean
Tr1 2 1 2 3 2
Tr2 4 4 1 1 2.5
Tr3 3 4 3 2 3
Block Mean 2 3 3 2 overall mean = 2.5

Consider the randomized block design with 4 blocks and 3 treatments given above. What is the total sum of squares?
Answer #1
Treatment 1 2 3 4 treatment mean
tr1 2 1 2 3 2
tr2 4 4 1 1 2.5
tr3 3 4 3 2 3
block mean 3 3 2 2

overall mean ==2.5

# the total sum of squares=SST

SSt=(Xij-)^2=(2-2.5)^2+(1-2.5)^2+(2-2.5)^2+(3-2.5)^2+(4-2.5)^2+(4-2.5)^2+(1-2.5)^2++(1-2.5)^2+(3-2.5)^2+(4-2.5)^2+(3-2.5)^2+(2-2.5)^2=15

# the total sum of squares=15